OpenStudy (anonymous):

How do I convert r=10cos(theta) to rectangular coordinates

OpenStudy (asnaseer):

what do you know about how rectangular and polar coordinates are related?

OpenStudy (asnaseer):

if you need to brush up on it then this might be of help: http://www.mathsisfun.com/polar-cartesian-coordinates.html let me know if you are still stuck after this

OpenStudy (anonymous):

I have an equation, but I don't have theta. I know that graphically r=10cos(theta) is a circle with a radius of 5, and the center of (5,0) but I have no idea how to find the rectangular coordinates. I know the equation of the circle is (x-5)^2+y^2=25 but I don't know if that is what the question is specifically asking for or not.

OpenStudy (asnaseer):

did you look at the link I gave above?

OpenStudy (anonymous):

Yes I did and I am having troubles figuring out what to do. I know that (x-5)^2+y^2=25, and I know that r = x^2 +y^2, and I know that tan^-1(Y/X) = Theta, but I need some help putting this all together to convert the equation r=10cos(theta) to rectangular coordinates

OpenStudy (asnaseer):

you mean \(r^2=x^2+y^2\)

OpenStudy (asnaseer):

you are given:\[r=10\cos(\theta)\]multiply both sides by 'r' to get:\[r^2=10r\cos(\theta)\]can you see how to proceed from here?

OpenStudy (anonymous):

It looks like we are trying to set up the equation to look like this: \[r ^{2} = x ^{2} + y ^{2}\] but I don't see what exactly can be done still because x and y is every single coordinate on the circle

OpenStudy (asnaseer):

no - that was in response to you stating "... I know that r = x^2 +y^2, ..."

OpenStudy (asnaseer):

the relation is not r = x^2 + y^2 its r^2 = x^2 + y^2

OpenStudy (asnaseer):

look at the equation I wrote above:\[r^2=10r\cos(\theta)\]notice that this can be written as:\[r^2=10(r\cos(\theta))\]now replace \(r^2\) and \(r\cos(\theta)\) with the rectangular equivalents

OpenStudy (anonymous):

That was a typo. I meant \[r ^{2} = x ^{2} + y ^{2}\]

OpenStudy (anonymous):

Which incidentally, that is also the equation of a circle...

OpenStudy (asnaseer):

polar and rectangular coordinates are related as follows:\[r^2=x^2+y^2\]\[x=r\cos(\theta)\]\[y=r\sin(\theta)\]\[\tan(\theta)=\frac{y}{x}\] make use of this to convert this equation to rectangular coordinates:\[r^2=10r\cos(\theta)\]

OpenStudy (anonymous):

So if I know the equation of the circle is \[(x-5)^{2} +y ^{2}=25\] and I know that \[r ^{2}=10(rcos(\theta))\] I can say that \[\tan ^{-1}(\theta) = \tan ^{-1}(0/-5) = 0 \] so then \[r^{2}=10(rcos(0))\]

OpenStudy (anonymous):

Am I on the right track?

OpenStudy (asnaseer):

no

OpenStudy (asnaseer):

you are stating off from a polar equation:\[r=10\cos(\theta)\]we then multiply both sides by 'r' to get:\[r^2=10r\cos(\theta)\]we then make use of the relationships between polar and rectangular coordinates to get:\[x^2+y^2=10x\]this can then be re-arranged into the standard equation for a circle

OpenStudy (asnaseer):